This is a quesion of two step equations.
The perimeter of a triangle is 51cm.The lenghts of its sides are sonsecutive odd integers.find the lenghts of all three sides.

Answer by Earlsdon(6294)

Let x be the first odd number, then x+2 is the next consecutive odd number and x+4 is the next one after that. so...
x+(x+2)+(x+4) = 51 Combine like-terms on the left.
3x+6 = 51 Subtract 6 from both sides.
3x = 45 Divide both sides by 3.
x = 15, x+2 = 17, and x+4 = 19
The lengths of the three sides are:
15cm, 17cm, and 19cm.